# Scalable Private Search with Wally

### v2

@@ 3 System Design
  • Hiding the content of real and fake queries (somewhat homomorphic encryption).
  • Removing the identifiable information from each query and shuffling the batch of queries received within a second (anonymization network).
@@
Within each epoch, Wally performs a semantic search that outputs the most relevant results to client queries, similar to insecure search engines.
@@
Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, it is used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [DCLT19].
In the following sections, we introduce the insecure version of Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
@@ 3.1 Insecure Search Pipeline
Once the client query and the server index are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
  • Server initialization. At initialization, the server divides the entries into $K$ clusters using $K$-means clustering. A heuristic-based technique partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$ and outputs these clusters with their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
  • Pick nearest centroid. To make a query, the client first finds the cluster nearest to the query embedding. For this, the client computes a cosine similarity between the query and each centroid embedding and picks a cluster whose centroid has high cosine similarity.
  • Server computation. The client then sends the nearest cluster ID $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes cosine similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
@@
The above protocol is not private. To ensure that the protocol does not reveal any information about the client, it must hide the following information from the server:
@@
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about client queries.
@@
3.2 Hiding the Query Embedding with SHE
To hide a query embedding, we rely on vectorized, RLWE-based SHE introduced in Section 2. The client query is an encrypted vector, and the server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
Server database.
The server divides the database, of size $N$, into $K$ clusters each with at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d*d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that each diagonal of a slice is a separate plaintext vector of size $n=d$. That is each slice is encoded as $d$ plaintexts vectors and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
Dot product computation.
Recall, from Section 2, for cosine similarity, the server needs to compute a dot-product between two embeddings. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $d=4$. This diagonal method is to minimize homomorphic rotations [HS14, HS18].
@@ 3.2 Hiding the Query Embedding with SHE
In the above description, we assume that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, the $j$-th plaintext vector in a cube holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
@@
Recall from Section 2 that homomorphic rotation is a costly operation. Even after packing, the server still performs $d$ rotations. Looking ahead to Section 6, we explain how we reduce the number of rotations to $2\sqrt{d}$ by using the standard the baby-step giant-step optimization [HS18].
@@
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. Wally converts each float into an integer by multiplying with the appropriate scaling factor and only taking the integer part of the result. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
@@
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding nearest cluster is a significant performance bottleneck in state-of-the-art schemes like Tiptoe. Specifically, to hide the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database. Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the strongest possible sense, but this results in a prohibitively high server computation.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is that the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic, obliviousness Tiptoe achieves.
The starting point of our solution is that in large-scale systems, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism allowing clients to hide their queries among the crowd of all queries. Our proposed approach is based on $(\epsilon,\delta)$-differential privacy. The approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
Intuitively, one would expect that routing all the traffic through an anonymization network would provide sufficient privacy as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can potentially learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known to not access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
Our protocol randomizes queries to guarantee differential privacy. Specifically, we make the following two changes:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks on of the clusters.
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into a slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
In Section 5, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client might want to retrieve metadata associated with it. When each entry has a small amount of metadata, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally picks the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run PIR to fetch the metadata of only the most relevant entry.
3.5 Keyword Search
Wally also supports keyword search in which the client is interested in retrieving data associated with a keyword. Using a hash function, the server database is divided into random $K^{\prime}$ buckets (instead of clusters). For each query, the client locally finds the bucket in which the keyword falls and then uses the keyword PIR scheme to fetch a specific record within a bucket.
@@ 3.5 Keyword Search
Wally uses a keyword PIR construction based on cuckoo hashing and index PIR. At a high level, at initialization, the server uses the cuckoo hashing to map the $N$ database elements into a table of size $O(N)$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [ALP+21] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
@@
Each bucket consisting of $N/K^{\prime}$ entries represented as $\sqrt{N/K^{\prime}}\times\sqrt{N/K^{\prime}}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
@@
    First expand $\hat{q}$ to two $\sqrt{N/K^{\prime}}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. To expand, the server uses the oblivious expansion algorithm given in [ALP+21]. This algorithm requires $2\sqrt{N/K^{\prime}}$ homomorphic substitutions.
@@
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/K^{\prime}$ plaintext-ciphertext multiplications and $N/K^{\prime}-\sqrt{N/K^{\prime}}$ ciphertext additions.
@@
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/K^{\prime}}$ ciphertext-ciphertext multiplications and $\sqrt{N/K^{\prime}}-1$ ciphertext additions. to compute.
The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule. Section 6 discusses various optimizations we implemented for MulPIR. Overall, our optimizations result in around a three times reduction in server computation and a two-three times reduction in request size.
[…]
Figure 4: Simplified protocol for Wally. Each client generates queries with Algorithm 2 and sends them to the anonymization network, $\mathcal{AN}$, according to the randomized schedule described in Algorithm 4. Cluster $j$ receives anonymous queries denoted as $(\tilde{q}^{j}_{i})$ and computes each response $(r^{j}_{i})$ via SHE.

### v3

@@ 3 System Overview
  • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
@@
Similar to state of the art insecure semantic search and previous secure search, Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, it is used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [15].
In the following sections, we introduce the insecure version of semantic search in Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
@@ 3.1 Insecure Search Pipeline
In the offline phase, the server maps each document to the corresponding embedding. At the time of a query the client embeds the query into an embedding. Once the client query and the server documents are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
  • Server initialization. At initialization, the server divides the documents (using their embeddings) into $K$ clusters using K-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
  • Pick nearest centroid. To make a query, the client first locally finds the cluster nearest to the query embedding. For this, the client computes a similarity between the query and each of the centroid embeddings and picks a cluster whose centroid has high similarity.
  • Server computation. The client then sends the nearest cluster index $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
@@
The above protocol is not private. To ensure that the protocol does not reveal any information about the client’s query, it must hide the following information from the server:
@@
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
@@
3.2 Hiding the Query Embedding
We rely on efficient vectorized, RLWE-based SHE described in Section 2 to hide a query embedding. At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster. The server then returns encrypted scores to the client. SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient. Overall, in our implementation, the client query and server response are a single RLWE ciphertext of sizes 226kB, and 53kB, respectively, (the former with evaluation keys) and it only takes 19 ms for the server to compute similarity scores for a single cluster homomorphically.
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding the nearest cluster is a significant performance bottleneck in previous schemes like Tiptoe. Specifically, to conceal the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database (all the clusters). Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the most robust possible sense, but this results in a prohibitively high server computation cost.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores only against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is a reduction in request size; concretely, the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic obliviousness Tiptoe achieves.
The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism to hide a particular client query among the crowd of queries from all the clients. Our proposed approach provides $(\epsilon,\delta)$-differential privacy. This approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
As the first step, we require all the traffic (queries and responses) to and from the server to route through the anonymization network. Intuitively, one would expect that doing this would hide the query in the crowd of queries, as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster then by observing that the cluster is accessed, the server can infer whether a request is from the client or not. Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client. For example, the queries towards each epoch’s start could belong to a particular user.
To achieve the desired privacy, we make the following two additional changes, these changes are highlighted in Figure 3:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks one of the clusters.
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
Intuitively, these two changes suppress the leakage, as mentioned above. In Section 6, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client retrieves metadata associated with it. When metadata per entry is small, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally selects the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run Keyword PIR to fetch the metadata of only the most relevant entry within a cluster.
3.5 Keyword Search
Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword. Using a random hash function, the server divides the database into random $K^{\prime}$ buckets (instead of clusters). For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket. The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule.
4 Efficient SHE-Based Instantiation
In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, the an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls. In this section, we discuss the details of our efficient SHE-based instantiation.
4.1 SHE details in semantic search
The server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
Server database.
Recall that the server database is divided into K clusters. Each cluster consists of at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d*d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that the plaintext dimension size $n=d$ and each diagonal within a slice is a separate plaintext. That is each slice is represented as $d$ plaintexts vectors, and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
@@ 4.1 SHE details in semantic search
SHE dot-product computation.
Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $n=d=4$.
@@
The above description assumes that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, a cube’s $j$-th plaintext vector holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
@@
Recall from Section 2 that homomorphic rotation is costly. In the above explanation, the server performs $d$ rotations even after packing to align the query. In Wally  we use the standard baby-step giant-step (BSGS) optimization [27] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step). See Algorithm 1 in [27] for the BSGS algorithm.
@@
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. In Wally we convert each float into an integer by multiplying it with the appropriate scaling factor having the plaintext modulus be large enough to perform the computation over the integers. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
@@
4.2 SHE details in keyword search
The server database consists of $N$ (keyword, value) entries. Recall that the server divides these entries into $B$ buckets using a random hash function. Within each bucket, in Wally  we use a keyword PIR construction based on cuckoo hashing and index PIR.
@@ 4.2 SHE details in keyword search
At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
@@
Each server bucket consisting of $N/B$ entries represented as $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
@@
    First expand $\hat{q}$ to two $\sqrt{N/B}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. The server uses the oblivious expansion algorithm given in [4] to expand. This algorithm requires $2\sqrt{N/B}$ homomorphic substitutions.
@@
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/B$ plaintext-ciphertext multiplications and $N/B-\sqrt{N/B}$ ciphertext additions.
@@
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/B}$ ciphertext-ciphertext multiplications and $\sqrt{N/B}-1$ ciphertext additions. to compute.
Section B discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.